OpenStudy (anonymous):

Simplify x^2-12x+20/x-10 A. x-2 B. x-2/x-10 C. x-2/10 D. x-10/x-2

OpenStudy (turingtest):

\[{x^2-12x+20 \over x-10}={(x-2)(x-10)\over x-10}=x-2\]